OpenStudy (anonymous):

4.One of your friends sends you an email asking you to explain how all of the following expressions have the same answer. 3 sqwrt x^3 x^1/3*x^1/3*x^1/3 1/x^-1 11sqwrt x^5*x^4*x^2 Compose an email back assisting your friend and highlight the names of the properties of exponents when you use them.

OpenStudy (anonymous):

\[3\sqrt{x ^{3}} \] \[x \frac{ 1 }{ 3 }*x \frac{ 1 }{ 3 }*x \frac{ 1 }{ 3}\] \[\frac{ 1 }{ x^{-1} }\] \[11\sqrt{x ^{5}}*x ^{4}*x ^{2}\]

OpenStudy (anonymous):

Compose an email back assisting your friend and highlight the names of the properties of exponents when you use them.

OpenStudy (anonymous):

\[x^(2/3)\] \[x^(1/3+1/3+1/3)=x^1\] \[1/x^-1=x^1\] \[x^((5/11)+4+2)=x^((5+44+22)/11)=x^(71/11)\]

OpenStudy (anonymous):

sorry first one is x^1

OpenStudy (anonymous):

Wow. Thank you so much!

OpenStudy (anonymous):

but 4th one is not x^1

OpenStudy (anonymous):

hmm that's werid... I might have made a mistake it looks like this: \[^{11}\]

OpenStudy (anonymous):

like little 11 then squrt, not big 11, if that makes scense

OpenStudy (anonymous):

What do they mean when they say highlight the names of the properties of exponents

OpenStudy (anonymous):

?

OpenStudy (anonymous):

its \[\sqrt[11]{x^5*x^4*x^2}=x^1\]

OpenStudy (anonymous):

oh cool! so I messed up lol...What do they mean when they say highlight the names of the properties of exponents?

OpenStudy (anonymous):

Yeah!!! Thank you for all your help ^-^

OpenStudy (anonymous):

Welcome